# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

### v1

@@ 7 Conclusion
In this paper, we introduce MiLoRA, a simple yet effective low-rank adaption method for LLMs. MiLoRA effectively learns on finetuning tasks while better preserving the pretrained knowledge by adapting the minor singular components of pretrained weight matrices. We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following. Experiment results demonstrate that MiLoRA consistently outperforms LoRA and PiSSA without sacrificing training or inference efficiency. We hope that our work will inspire future research on parameter-efficient finetuning of LLMs.
@@ Appendix A Detailed Experiment Setups
A.1 Similarity metric between subspace
@@
A.2 Our hyperparameter setup
Table 8 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023b), therefore we denote it as LLM-adapters setup.
@@ A.2 Our hyperparameter setup
A.3 PiSSA hyperparameter setup
@@
Table 9: Hyperparameter configuration from the PiSSA paper.
@@
A.4 Computation resource
All of our experiments are conducted on 8 NVIDIA L40 GPUs.
@@ Appendix B More Experiment Results
B.1 Detailed results on math reasoning across ranks
Detailed EM scores of MiLoRA and baselines on math reasoning by adjusting rank $r$ within the set $\{16,32,64,128\}$.
Table 10: Math reasoning evaluation results for LLaMA2-7B across different ranks.
Method Dataset 16 32 64 128
LoRA GSM8K 62.93 63.38 63.61 59.59
MATH 15.52 15.56 15.28 15.28
PiSSA GSM8K 61.56 60.12 60.65 56.94
MATH 13.00 14.36 14.60 14.76
MiLoRA GSM8K 62.55 63.53 63.99 61.56
MATH 14.12 15.06 16.14 16.16
B.2 More results of the relationship between $\Delta W$ and $W$
@@
B.3 Math reasoning evaluation results for LLaMA2-7B under PiSSA configuration
Table 12 presents math reasoning evaluation results for LLaMA2-7B finetuned for 10 epoches under PiSSA configuration. The best checkpoint is selected on a validation set.
Table 12: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH finetuned 10 epoches under PiSSA configuration.
Method GSM8K MATH Avg.
LoRA 56.71 10.28 33.50
PiSSA 55.72 10.02 32.87
MiLoRA 55.57 9.08 32.33
Appendix C Limitations and Broader Impacts
Limitations
Due to limited computational resources, following previous work, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
Broader Impacts

### v2

@@ 7 Conclusion
In this paper, we introduce MiLoRA, a simple yet effective low-rank adaption method for LLMs. MiLoRA effectively learns on finetuning tasks while better preserving the pretrained knowledge by adapting the minor singular components of pretrained weight matrices. We investigate the effectiveness of MiLoRA on a wide range of LLM evaluation benchmarks, including commonsense reasoning, math reasoning and instruction-following, and vision-language model evaluation benchmarks. Experiment results demonstrate that MiLoRA consistently outperforms LoRA and PiSSA without sacrificing training or inference efficiency. We hope that our work will inspire future research on parameter-efficient finetuning of LLMs.
Limitations
Due to limited computational resources, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks following previous work Liu et al. (2024b); Wu et al. (2024b). We examine and compare MiLoRA with baselines using the hyperparameter configurations provided in the previous works Hu et al. (2023); Liu et al. (2024b) instead of exhaustive hyperparameter search for each task. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
@@ Appendix A Detailed Experiment Setups
A.1 Similarity Metric Between Subspaces
@@
A.2 Our Hyperparameter Setup for LLM
Table 10 shows our detailed hyperparameters. This setup follows LLM-adapters (Hu et al., 2023), therefore we denote it as LLM-adapters setup.
@@ A.2 Our Hyperparameter Setup for LLM
Visual instruction tuning hyperparameters
Rank r 128
$\alpha$ of LoRA 256
$\alpha$ of PiSSA/MiLoRA 128
Dropout 0.05
Optimizer AdamW
LR 2e-4
LR Scheduler cosine
Batch Size 16
Warmup Ratio 0.03
Epochs 1
Placement query, key, value, output, gate, MLP up, MLP down
Table 11: Our hyperparameter configurations applied in visual instruction tuning tasks.
A.3 PiSSA Hyperparameter Setup for LLM
@@
Table 12: The detailed hyperparameter configurations used by PiSSAMeng et al. (2024).
A.4 More Implementation Details
We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning. For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Table 11 shows the details.
@@ Appendix B More Experiment Results
Appendix C Broader Impacts